E-Z Stop Fast Gas sold $10,957 worth of gasoline yesterday. Regular sold for $2.30 a gallon, and premium sold for $2.55 a gallon. If the station sold 420 more gallons of regular than of premium: How many premium gallons / How many Regular Gallons?
I don't know quite how to put this equation together. Please someone help and explain it to me. Please Please Please

Answer by umnikmo(75)

Lets say the station sold G gallons of regular, then it also sold G-420 gallons of premium. now we can set up an equation:
2.3G+2.55(G-420)=10957
2.3G+2.55G-1071=10957
4.85G=9886
G=2480 - gallons of regular sold
G-420=2060 - gallons of premium sold